# Mitigating Forgetting in LLM Fine-Tuning via Low-Perplexity Token Learning

### v3

@@ Abstract
Maintaining consistent model performance across domains is a fundamental challenge in machine learning. While recent work has explored using LLM-generated data for fine-tuning, its impact on cross-domain generalization remains poorly understood. This paper presents a systematic analysis revealing that fine-tuning with LLM-generated data not only improves target task performance but also reduces non-target task degradation compared to fine-tuning with ground truth data. Through analyzing the data sequence in tasks of various domains, we demonstrate that this enhancement of non-target task robustness stems from the reduction of high perplexity tokens found in LLM-generated sequences. Following our findings, we showed that masking high perplexity tokens in ground truth training data achieves similar non-target task performance preservation, comparable to using LLM-generated data. Extensive experiments across different model families and scales, including Gemma 2 IT 2B, Llama 3 8B Instruct, and 3 additional models, agree with our findings. To the best of our knowledge, this is the first work to provide an empirical explanation based on token perplexity reduction to mitigate catastrophic forgetting in LLMs after fine-tuning, offering valuable insights for developing more robust fine-tuning strategies.
@@ 1 Introduction
Supervised fine-tuning (SFT) of large language models (LLMs) [26] has proven highly effective for enhancing their ability to follow novel instructions [37] and produce useful outputs across a wide range of tasks, including summarization and web querying [31, 22, 1, 16]. As LLMs are increasingly applied in specialized domains such as arithmetic and programming assistance [33, 23], practitioners often encounter substantial computational and data cleaning issues when adapting these models to gain improvements. These difficulties are exacerbated by performance saturation in smaller LLMs, where benchmark improvements on tasks like MATH [12] and MBPP [3] tend to plateau as model size decreases. Moreover, fine-tuning can risk degrading a model’s general capabilities [28, 20, 27, 4]. This raises a critical question: how can we efficiently fine-tune pre-trained, instruction-following LLMs for domain-specific tasks without compromising other skill sets such as arithmetic or common sense reasoning, especially when the original fine-tuning data is unavailable or computational resources are constrained?
Recent work has shown that fine-tuning with model-generated data can be highly effective, often outperforming training on ground truth data when using high-quality generated samples [42, 10, 29]. For example, Ren et al. [28] demonstrated that model-generated responses can surpass human annotations in downstream performance, particularly when using more capable models (e.g., GPT-4) to generate training data and mitigate degradation on general tasks. Although they attributed part of this success to the “familiarity” of generated responses, measured via low perplexity, it remains unclear why such data works well for instruction tuning. Specifically: (1) The regenerated response may differ contextually from the ground truth, and such contextual bias could affect the utility of low-perplexity training. (2) The “non-forgetting” impact of low-perplexity training on the performance of tasks not being fine-tuned, which maintains the practicality of the model, is still underexplored. (3) Since model size plays a role in distillation benefits, it is important to disentangle the effects of model scale and evaluate low-perplexity training using models of the same size.
@@
Self-Output (Llama 3 8B Instruct): Let $r_{1}$ and $r_{2}$ be the roots of $45x^{2}+3x+$ 4. Then, by Vieta’s formulas, we have $r_{1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+r}_{2}=-\frac{3}{5}$ and $r_{1}r_{2}=\frac{4}{5}$. Let $\alpha=\frac{1}{r_{1}}$ and $\beta=\frac{1}{r_{2}}$. Then, we have $\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+}\beta=\frac{1}{r_{1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+}\frac{1}{r_{2}}=\frac{r_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2}}+r_{1}}{r_{1}r_{2}}=\frac{-\frac{3}{5}}{\frac{4}{5}}=\boxed{-\frac{3}{4}}$
@@ 2.1 Training and Evaluation Framework
Table 1: Perplexity average and variance over answer sequences of MBPP and MATH training datasets on Llama 3-8B Instruct. The average sentence perplexity is calculated by averaging the sum of each sentence’s perplexity over the token perplexity score.
@@ 2.2 Self-Output
Self-Output follows a similar high-level approach, similar to other synthetic data generation techniques [42, 28], where a high-quality synthetic dataset is created by sampling diverse responses from an LLM and applying a strict filtering process. For each training instance, we use a language model $M$ (e.g., Llama 3 8B Instruct) to generate $N$ (e.g., 32) distinct responses with a temperature setting of $T=0.7$. The generated responses are then filtered to retain only those that are semantically aligned with the ground truth, ensuring the quality and relevance of the synthetic data.
@@ 2.3 Rephrase
Given $T$ tasks composed of one target task and $T-1$ non-target tasks, TI is the target task performance improvement percentage after training on the target task. BWT refers to the average performance degradation percentage of the $T-1$ non-target task after training on the target task. The models to fine-tune in this paper refer to generic instruction following models before fine-tuning with target task data, e.g., Llama 3 8B Instruct [8], Gemma 2 IT 2B [32], Mistral 7B Instruct [15], and OLMo 2 7B series [24]. For target task performance, Self-Output demonstrates consistent positive TI in both MBPP and MATH. For the non-target task, the Self-Output method achieves better BWT scores than the baseline fine-tuning in most cases, indicating much less degradation on non-target task capabilities. For more details, such as performance comparison with other models or other minor findings for self-output and rephrase method, please refer to Appendix B.
@@ 2.4 Analysis of perplexity across each sample
In this section, we analyze the perplexity of datasets generated by Self-Output and Rephrase to quantify their impact on model uncertainty and fine-tuning stability. We first collect the inference statistics of GPT-4o, GPT-4o mini, Gemma 2 IT 27B, Llama 3 8B Instruct, and Gemma 2 IT 2B from all three methods: Baseline Fine-tuning, Rephrase, and Self-Output strategy, along with their BWT, as Figure 3 shows that a higher perplexity of target training data could result in a higher degradation on non-target task. Second, we further collect the Llama 3 8B Instruct inference statistics from three kinds of training data: Baseline Finetuning (ground truth data), Rephrase, and Self-Output strategy. We evaluate the perplexity of responses by summing up the negative log probability of each token conditioned on its previous context in the sentence to measure the uncertainty of model outputs over the entire sequence. Specifically, the perplexity $PPL$ [^1] In this section, we use natural log instead of base 2 for computation convenience across a sequence is computed as:
@@
Table 1 demonstrates that the Self-Output strategy achieves the lowest average perplexity of 1.16 in MBPP and 1.34 in MATH. Among the three approaches, Ground Truth exhibits the highest uncertainty with a mean perplexity of 4.83, and the perplexity of Rephrase is slightly higher than Self-Output. The lower perplexity of Self-Output suggests that LLMs are less likely to deviate far from their initial weights when fine-tuning on downstream tasks created by Self-Output. This characteristic maintains the model’s original performance on non-target tasks while effectively learning from the target task training data.
To investigate predictability at the token-level, we sample a single MATH response and visualize the token-level perplexity value. Figure 3 reveals distinct patterns across Ground Truth, Rephrase, and Self-Output responses. The first 20 tokens’ variance is 2.04 for Ground Truth, 2.31 for Rephrase, and 0.04 for Self-Output, which confirms our observation: Ground Truth exhibits frequent perplexity spikes of high amplitude (perplexity around 6-8), which reflects the natural variability in human problem-solving expressions. While Rephrase maintains similar peak amplitudes to Ground Truth, i.e., more regular spike patterns, its variability is between Ground Truth and Self-Outputs. Self-Output demonstrates remarkably consistent low token perplexity values rarely exceeding 2, indicating highly predictable token sequences. These patterns suggest that Self-Output generates more deterministic solutions, potentially beneficial for maintaining consistent problem-solving approaches, while Ground Truth sequences introduce much higher variations. To demonstrate the different content of tokens distribution between Ground Truth and Self-Output responses, please refer to Appendix C.
@@ 3 Low-perplexity token learning via selective token masking
Our earlier analysis (Figure 3) reveals a notable discrepancy in token perplexity distributions between Self-Output generated data and other sources. This observation raises a central question: Is the superior performance of Self-Output methods primarily due to the lower perplexity of their generated tokens? To explore this hypothesis and design more efficient fine-tuning strategies, we propose Selective Token Masking (STM), a novel approach to supervised fine-tuning (SFT) that explicitly leverages token-level perplexity.
@@
While our approach shares conceptual similarities with prior work such as Lin et al. [19], Ankner et al. [2], it differs significantly in both implementation and efficiency. Previous methods typically involved a two-stage process: training a reference model on high-quality data to identify learnable (high-perplexity) tokens, followed by pre-training a larger LLM on this curated subset. In contrast, STM achieves similar goals in a streamlined single-stage workflow by directly using an existing instruction-tuned model to compute token perplexities in a single forward pass. This design avoids the need to train or maintain any additional models, significantly reducing computational overhead while preserving the benefits of low-perplexity training. This makes STM a compelling choice for efficient fine-tuning rather than full-scale pretraining.
@@ 4 Experimental Setting
To evaluate the effectiveness of our proposed method, we adopt benchmark datasets specifically selected to assess model robustness on both target and non-target tasks after fine-tuning. We include only datasets with verifiable ground truth, allowing us to filter out incorrect self-generated responses for a fair comparison with self-synthesized baselines and to ensure high data quality. Following recent findings [4] that LoRA [13] effectively mitigates performance degradation, we standardize our experiments by using LoRA for all fine-tuning. We then vary training strategies and data sources to study their impact on performance preservation. Details of our evaluation prompts and training configurations are provided in Appendix D. All implementation code and datasets are publicly available at: https://github.com/appier-research/robust-llm-finetunes.
@@
4.1 Target and non-target datasets
We focus on three domains, programming, mathematics, and knowledge-based tasks, to study performance degradation after supervised fine-tuning with LoRA on each domain. For the programming and mathematics domains, we select MBPP and MATH for target task training, while others and knowledge-based tasks for non-target task evaluation to assess generalization capabilities.
4.1.1 Programming Tasks
For code generation evaluation, we use MBPP for both training and evaluation, while BIRD serves as an out-of-domain assessment. This pairing tests both direct programming ability and cross-language generalization from Python to SQL.
4.1.2 Mathematical Reasoning Tasks
We use MATH for training and evaluation, and GSM8K for out-of-domain testing. These datasets differ in format: MATH features competition-style problems, while GSM8K uses natural language, allowing us to assess generalization across mathematical expression styles.
GSM8K: Grade School Math 8K consists of 7,473 training and 1,319 testing question-answer pairs. We utilize only the test set for out-of-domain evaluation, leveraging its natural language format to assess generalization from formal to informal mathematical reasoning.
4.1.3 Knowledge-based Task
To assess broader generalization capabilities beyond mathematical and programming domains, we incorporate:
ARC-Challenge: A specialized subset of the ARC science question dataset, comprising 2,590 questions (1,119 training, 299 validation, and 1,172 test) selected for their increased difficulty and reduced susceptibility to statistical shortcuts. This dataset is used for evaluation.
4.2 Models
These models were selected to represent different points in the compute-performance trade-off spectrum. Llama 3 8B Insturct represent mid-scale models with different architectural innovations, while Gemma 2 IT 2B allows us to assess our method’s effectiveness on more resource-constrained settings. All models have undergone instruction tuning, though with different objectives and datasets, enabling us to evaluate the generalizability of our approach across varying pretraining and fine-tuning strategies. To further test the generalization of STM effectiveness on a brand new model we also choose three additional new models to test on target task MBPP and MATH, please refer to Appendix B.
@@ 5 STM Results
Table 2 shows that with a threshold to filter out high perplexity tokens such as 2.5, which approximately filters out around 20% to 24% of total tokens, its BWT is close to $0\%$, and STM also achieves comparable TI to the Self-Output method. It indicates that STM not only improves target task performance but also reduces non-target task degradation. This result agrees with our hypothesis that the presence of high perplexity tokens is one of the causes of performance degradation of the ground truth model, and low perplexity tokens sufficiently mitigate such negative impact for strong generalization without the need for self-generated data. We also applied STM to other models, and the results can be found in Appendix B. Our conclusions are still the same across different models.
@@ 5.1 Does filtering high perplexity tokens always perform better?
From the intuition of the high perplexity filtering mechanism of STM, we have done an ablation study on the masking criteria of perplexity by filtering out high, random, and low ppl tokens. More specifically, high ppl refers to filtering the top 25% ppl of all tokens, random filtering refers to filtering the randomly selected 25% of all tokens, and low ppl refers to filtering the lowest 25% ppl of all tokens from the dataset. As the upper part of Table 4 and Table 4 show, masking high ppl tokens in STM simultaneously reduces performance degradation and improves target task performance. Likewise, we have the same conclusions for Llama 3 8B Instruct on such ablation test, please refer to Appendix B.
Table 3: Ablation and scaling on threshold of STM with Gemma 2 IT 2B on MBPP task. Percentages indicate the ratio of tokens filtered from training data. STM performs generally the best with 20% to 24% of high ppl tokens filtered.
BWT(%) TI(%)
@@
Table 4: Ablation and scaling on threshold of STM with Llama3 8B Instruct on MBPP and MATH task respectively. Percentages indicate the ratio of tokens filtered from training data. STM performs generally the best with 20% to 24% of high ppl tokens filtered.
BWT(%) TI(%)
STMτ=2.5,high   0.16   3.12
STMτ=2.5,random -6.64 -9.84
STMτ=2.5,low -9.15 -14.73
$STM_{\tau=1000}$ (1.7%) -19.6 -11.8
$STM_{\tau=10}$ (11.0%) -13.1   5.1
$STM_{\tau=2.5}$ (22.2%)   0.16   3.12

### v6

@@ Abstract
Maintaining consistent model performance across domains is a fundamental challenge in machine learning. While recent work has explored using LLM-generated data for fine-tuning, its impact on cross-domain generalization remains poorly understood. This paper presents a systematic analysis revealing that fine-tuning with LLM-generated data not only improves target task performance but also reduces non-target task degradation compared to fine-tuning with ground truth data. Through analyzing the data sequence in tasks of various domains, we demonstrate that this enhancement of non-target task robustness stems from the reduction of high perplexity tokens found in LLM-generated sequences. Following our findings, we showed that masking high perplexity tokens in ground truth training data achieves similar non-target task performance preservation, comparable to using LLM-generated data. Extensive experiments across different model families and scales, including Gemma 2 IT 2B, Llama 3 8B Instruct, and three additional models, agree with our findings. To the best of our knowledge, this is the first work to provide an empirical explanation based on token perplexity reduction to mitigate catastrophic forgetting in LLMs after fine-tuning, offering valuable insights for developing more robust fine-tuning strategies. footnotetext: Our code and datasets are available at https://github.com/appier-research/robust-llm-finetunes
@@ 1 Introduction
Supervised fine-tuning large language models (LLMs) [29] has proven highly effective for enhancing their ability to follow novel instructions [43, 21, 23, 15, 49] and produce useful outputs across a wide range of tasks, including summarization and web querying [34, 26, 1, 17]. As LLMs are increasingly applied in specialized domains such as arithmetic and programming assistance [38, 27], practitioners often encounter substantial computational and data cleaning issues when adapting these models to gain improvements. These difficulties are exacerbated by performance saturation in smaller LLMs, where benchmark improvements on tasks like MATH [12] and MBPP [3] tend to plateau as model size decreases. Moreover, fine-tuning can risk degrading a model’s general capabilities [31, 24, 30, 4]. This raises a critical question: how can we efficiently fine-tune pre-trained, instruction-following LLMs for domain-specific tasks without compromising other skill sets such as arithmetic or common sense reasoning, especially when the original fine-tuning data is unavailable or computational resources are constrained?
Recent work has shown that fine-tuning with model-generated data can be highly effective, often outperforming training on ground truth data when using high-quality generated samples [48, 10, 32, 35]. For example, Ren et al. [31] demonstrated that model-generated responses can surpass human annotations in downstream performance, particularly when using more capable models (e.g., GPT-4) to generate training data and mitigate degradation on general tasks. Although they attributed part of this success to the “familiarity” of generated responses, measured via low perplexity, it remains unclear why such data works well for instruction tuning. Specifically: (1) The regenerated response may differ contextually from the ground truth, and such contextual bias could affect the utility of low-perplexity training. (2) The “non-forgetting” impact of low-perplexity training on the performance of tasks not being fine-tuned, which maintains the practicality of the model, is still underexplored. (3) Since model size plays a role in distillation benefits, it is important to disentangle the effects of model scale and evaluate low-perplexity training using models of the same size.
@@
Self-Output (Llama 3 8B Instruct): Let $r_{1}$ and $r_{2}$ be the roots of $5x^{2}+3x+4$. Then, by Vieta’s formulas, we have $r_{1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+r}_{2}=-\frac{3}{5}$ and $r_{1}r_{2}=\frac{4}{5}$. Let $\alpha=\frac{1}{r_{1}}$ and $\beta=\frac{1}{r_{2}}$. Then, we have $\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+}\beta=\frac{1}{r_{1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+}\frac{1}{r_{2}}=\frac{r_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2}}+r_{1}}{r_{1}r_{2}}=\frac{-\frac{3}{5}}{\frac{4}{5}}=\boxed{-\frac{3}{4}}$
@@ 2.1 Training and evaluation framework
Table 1: Perplexity average and variance over answer sequences of MBPP and MATH training datasets on Llama 3 8B Instruct. The average sentence perplexity is calculated by averaging the sum of each sentence’s perplexity over the token perplexity score.
@@ 2.2 Self-Output
Self-Output follows a similar high-level approach, similar to other synthetic data generation techniques [48, 31], where a high-quality synthetic dataset is created by sampling diverse responses from an LLM and applying a strict filtering process. For each training instance, we use a language model $M$ (e.g., Llama 3 8B Instruct) to generate $N$ (e.g., 32) distinct responses with a temperature setting of $T=0.7$. The generated responses are then filtered to those that are semantically aligned with the ground truth, ensuring the quality and relevance of synthetic data.
@@ 2.3 Rephrase
We have a sanity check on the output of Rephrase, and remove the incorrect label output as the final training set for Rephrase to ensure that all the final outputs of Rephrase are correct. But we did not evaluate the intermediate output of Rephrase, e.g., the reasoning or CoT process. This process is partially reflected in Appendix F, Table 17.
@@
Given $T$ tasks composed of one target task and T1T-1 non-target tasks, TI is the target task performance improvement percentage after training on the target task. BWT refers to the average performance degradation percentage of the T1T-1 non-target task after training on the target task. The models to fine-tune in this paper refer to generic instruction following models before fine-tuning with target task data, e.g., Llama 3 8B Instruct [8], Gemma 2 IT 2B [36], Mistral 7B Instruct [16], and OLMo 2 7B series [40]. For target task performance, Self-Output demonstrates consistent positive TI in both MBPP and MATH. For the non-target task, the Self-Output method achieves better BWT scores than the baseline fine-tuning in most cases, indicating much less degradation on non-target task capabilities. For more details, such as performance comparison with other models or other minor findings for self-output and rephrase method, please refer to Appendix D.
@@ 2.4 Analysis of perplexity across each sample
In this section, we analyze the perplexity of datasets generated by Self-Output and Rephrase to quantify their impact on model uncertainty and fine-tuning stability. We first collect the inference statistics of GPT-4o, GPT-4o mini, Gemma 2 IT 27B, Llama 3 8B Instruct, and Gemma 2 IT 2B from all three methods: Baseline Fine-tuning, Rephrase, and Self-Output strategy, along with their BWT, as Figure 3 shows that a higher perplexity of target training data could result in a higher degradation on non-target task. Second, we further collect the Llama 3 8B Instruct inference statistics from three kinds of training data: Baseline Finetuning (ground truth data), Rephrase, and Self-Output strategy. We evaluate the perplexity of responses by summing up the negative log probability of each token conditioned on its previous context in the sentence to measure the uncertainty of model outputs over the entire sequence. Specifically, the perplexity PPLPPL [^1] In this section, we use natural log instead of base 2 for computation convenience across a sequence is computed as:
@@
Table 1 demonstrates that the Self-Output strategy achieves the lowest average perplexity of 1.16 in MBPP and 1.34 in MATH. Among the three approaches, Ground Truth exhibits the highest uncertainty with a mean perplexity of 4.83, and the perplexity of Rephrase is slightly higher than Self-Output. The lower perplexity of Self-Output suggests that LLMs are less likely to deviate far from their initial weights when fine-tuning on downstream tasks created by Self-Output. This characteristic maintains the model’s original performance on non-target tasks while effectively learning from the target task.
To investigate predictability at the token-level, we sample a single MATH response and visualize the token-level perplexity value. Figure 3 reveals distinct patterns across Ground Truth, Rephrase, and Self-Output responses. The first 20 tokens’ variance is 2.04 for Ground Truth, 2.31 for Rephrase, and 0.04 for Self-Output, which confirms our observation: Ground Truth exhibits frequent perplexity spikes of high amplitude (perplexity around 6-8), which reflects the natural variability in human problem-solving expressions. While Rephrase maintains similar peak amplitudes to Ground Truth, i.e., more regular spike patterns, its variability is between Ground Truth and Self-Outputs. Self-Output demonstrates remarkably consistent low token perplexity values rarely exceeding 2, indicating highly predictable token sequences. These patterns suggest that Self-Output generates more deterministic solutions, potentially beneficial for maintaining consistent problem-solving approaches, while Ground Truth sequences introduce much higher variations. To demonstrate the different content of tokens distribution between Ground Truth and Self-Output responses, please refer to Appendix E.
@@ 3 Low-perplexity token learning via selective token masking
Our earlier analysis of Figure 3 reveals a notable discrepancy in token perplexity distributions between Self-Output generated data and other sources. This observation raises a central question: Is the superior performance of Self-Output methods primarily due to the lower perplexity of their generated tokens? To explore this hypothesis and design more efficient fine-tuning strategies, we propose Selective Token Masking (STM), a novel approach to supervised fine-tuning (SFT) that explicitly leverages token-level perplexity.
@@
While our approach shares conceptual similarities with prior work such as [20, 2], it differs significantly in both implementation and efficiency. Previous methods typically involved a two-stage process: training a reference model on high-quality data to identify learnable (high-perplexity) tokens, followed by pre-training a larger LLM on this curated subset. In contrast, STM achieves similar goals in a streamlined single-stage workflow by directly using an existing instruction-tuned model to compute token perplexities in a single forward pass. This design avoids the need to train or maintain any additional models, significantly reducing computational overhead while preserving the benefits of low-perplexity training. This makes STM a compelling choice for efficient fine-tuning rather than full-scale pretraining.
@@ 4 Experimental setting
To evaluate the effectiveness of our proposed method, we adopt benchmark datasets specifically selected to assess model robustness on both target and non-target tasks after fine-tuning. We include only datasets with verifiable ground truth, allowing us to filter out incorrect self-generated responses for a fair comparison with self-synthesized baselines and to ensure high data quality. Following recent findings [4] that LoRA [13] effectively mitigates performance degradation, we standardize our experiments by using LoRA for all fine-tuning. We then vary training strategies and data sources to study their impact on performance preservation. Details of our evaluation prompts and training configurations are provided in Appendix F. All implementation code and datasets are publicly available at: https://github.com/appier-research/robust-llm-finetunes.
4.1 Target and non-target datasets
We focus on three domains, programming, mathematics, and knowledge-based tasks, to study performance degradation after supervised fine-tuning with LoRA on each domain. For the programming and mathematics domains, we select MBPP and MATH for target task training, while others (MATH or MBPP test split, BIRD, GSM8K) and knowledge-based tasks for non-target task evaluation to assess generalization capabilities.
Programming. For code generation evaluation, we split MBPP into target task training data and non-target task testing data, while BIRD serves as a non-target assessment. This pairing tests both direct programming ability and cross-language generalization from Python to SQL.
Mathematical reasoning. We split MATH into target task training data and non-target task testing data, and GSM8K for non-target task testing. These datasets differ in format: MATH features competition-style problems, while GSM8K uses natural language, allowing us to assess generalization across mathematical expression styles.
Knowledge-based. To assess broader generalization capabilities beyond mathematical and programming domains, we incorporate ARC-Challenge, a subset of the ARC science question dataset.
For details about the data curation of the target and non-target datasets, please refer to Appendix A.
To explore more performances of non-target tasks including instruction following and safety performance under limited model choices, please refer to Appendix B.
@@ 5 STM results
Table 2 shows that with a threshold to filter out high perplexity tokens, such as 2.5, which approximately filters out around $20\%$ to $24\%$ of total tokens, its BWT is close to $0\%$, and STM also achieves comparable TI to the Self-Output method. It indicates that STM not only improves target task performance but also reduces non-target task degradation. This result agrees with our hypothesis that the presence of high perplexity tokens is one of the causes of performance degradation of the ground truth model, and low perplexity tokens sufficiently mitigate such negative impact for strong generalization without the need for self-generated data. We also applied STM to other models, and the results can be found in Appendix D. Our conclusions are still the same across different models.
@@ 5.1 Does filtering high perplexity tokens always perform better?
From the intuition of the high perplexity filtering mechanism of STM, we have done an ablation study on the masking criteria of perplexity by filtering out high, random, and low ppl tokens. More specifically, high ppl refers to filtering the top $25\%$ ppl of all tokens, random filtering refers to filtering the randomly selected $25\%$ of all tokens, and low ppl refers to filtering the lowest $25\%$ ppl of all tokens from the dataset. As the upper part of Table 4 shows, masking high ppl tokens in STM simultaneously reduces performance degradation and improves target task performance. Likewise, we have the same conclusions for Llama 3 8B Instruct on such ablation test, please refer to Appendix D.
Table 3: Ablation and scaling on threshold of STM with Gemma 2 IT 2B on MBPP task. Percentages indicate the ratio of tokens filtered from training data. STM performs generally the best with $20\%$ to $24\%$ of high ppl tokens filtered.
Configuration BWT(%) TI(%)